# PrivTune: Efficient and Privacy-Preserving Fine-Tuning of Large Language Models via Device-Cloud Collaboration

### v2

@@ V-D Token Importance-aware $d\chi$-Privacy
To provide formal privacy guarantees, we adopt the $d_{\chi}$}-Privacy mechanism [wang2024preference] instead of conventional DP. The $d_{\chi}$}-Privacy framework defines a smooth sensitivity bound tailored to continuous spaces such as embedding spaces, making it especially suitable for LLM representation perturbation.
@@
For classification tasks, we use the statistical features of tokens to represent their importance. We adopt a TF-IDF-inspired metric [aizawa2003information] to estimate the importance of token $x_{i}$ in a classification context. Let $p(x=x_{i}|y=c)$ be the frequency of token $x_{i}$ in class $c$, then the importance score ($IS$) of token $x_{i}$ is given by:
@@ VI-A Experiment Settings
The effectiveness of the PrivTune is evaluated on both text classification and generation tasks. For the classification tasks, we use the Financial Phrasebank (FP) [malo2014good] and Stanford Sentiment Treebank (SST) [wang2019glue], and the Blog Authorship Corpus [lyu2020differentially]. For the generation tasks, we use the question-answering datasets, i.e., TruthfulQA (TQA) [lin2022truthfulqa] and Fitness_Unformatted (FU) [hazsylvia_fitness_unformatted_2024] datasets.
@@
RoBERTa-Large [liu2019roberta], with 355 million parameters, Mistral-7B-v0.2 [jiang2023clip], with 7 billion parameters, and Llama-3 [llama3modelcard], with 8 billion parameters, are used as cloud models, both of which are publicly available at https://huggingface.co/.
@@
PrivTune is evaluated against four baseline methods: 1) DP-Forward [du2023dp], which directly perturbs the embedding matrix during the forward pass of the bottom model to meet strict DP requirements; 2) SLFT-DP [WANG2024Selective], which applies sequence-level DP to perturb embeddings in the forward pass for privacy-preserving fine-tuning within the SLFT framework; 3) SAP [shen2025sap], which integrates text privatization (i.e., $d_{\chi}$-Privacy) into an SL framework; 4) ScaleOT [yao2025scaleot], which employs an offsite-tuning framework to achieve privacy preservation; and 5) FedBiOT [wu2024fedbiot], which combines FL with offsite-tuning to enable secure and private fine-tuning.
@@
To comprehensively evaluate the privacy performance of the PrivTune, we employ two types of state-of-the-art privacy inference attacks, i.e., EIAs and AIAs attacks  [song2020information, zhu2019deep, shokri2017membership, du2023dp, huang2024transferable]. Specifically, we follow their technical implementations (i.e., original settings) and utilize their source codes to reproduce three representative EIA (i.e., Attack-0–Attack-2) and three representative AIA (i.e., Attack-3–Attack-5) attacks for evaluating PrivTune.
@@
We adopt the PEFT method, specifically LoRA [hulora], in PrivTune, and employ the AdamW optimizer with a linear learning rate scheduler during fine-tuning, with an initial learning rate of $3\times 10^{-4}$. By default, we set $l=3$, $n=3$, $k=2$, and $\delta=0.6$ in our experiments. For evaluation, we use classification accuracy and ROUGE-L score to measure the Utility Performance (UP) on classification and generation tasks, respectively [li2023privacy]. Additionally, Attack Success Rate (ASR) is used to assess the defense capability.

### v3

@@ V-D Token Importance-aware $d\chi$-Privacy
To provide formal privacy guarantees, we adopt the $d_{\chi}$}-Privacy mechanism [46] instead of conventional DP. The $d_{\chi}$}-Privacy framework defines a smooth sensitivity bound tailored to continuous spaces such as embedding spaces, making it especially suitable for LLM representation perturbation.
@@
For classification tasks, we use the statistical features of tokens to represent their importance. We adopt a TF-IDF-inspired metric [3] to estimate the importance of token $x_{i}$ in a classification context. Let $p(x=x_{i}|y=c)$ be the frequency of token $x_{i}$ in class $c$, then the importance score ($IS$) of token $x_{i}$ is given by:
@@ VI-A Experiment Settings
The effectiveness of the PrivTune is evaluated on both text classification and generation tasks. For the classification tasks, we use the Financial Phrasebank (FP) [25] and Stanford Sentiment Treebank (SST) [38], and the Blog Authorship Corpus [23]. For the generation tasks, we use the question-answering datasets, i.e., TruthfulQA (TQA) [19] and Fitness_Unformatted (FU) [12] datasets.
@@
RoBERTa-Large [21], with 355 million parameters, Mistral-7B-v0.2 [15], with 7 billion parameters, and Llama-3 [2], with 8 billion parameters, are used as cloud models, both of which are publicly available at https://huggingface.co/.
@@
PrivTune is evaluated against four baseline methods: 1) DP-Forward [7], which directly perturbs the embedding matrix during the forward pass of the bottom model to meet strict DP requirements; 2) SLFT-DP [40], which applies sequence-level DP to perturb embeddings in the forward pass for privacy-preserving fine-tuning within the SLFT framework; 3) SAP [32], which integrates text privatization (i.e., $d_{\chi}$-Privacy) into an SL framework; 4) ScaleOT [48], which employs an offsite-tuning framework to achieve privacy preservation; and 5) FedBiOT [43], which combines FL with offsite-tuning to enable secure and private fine-tuning.
@@
To comprehensively evaluate the privacy performance of the PrivTune, we employ two types of state-of-the-art privacy inference attacks, i.e., EIAs and AIAs attacks  [34, 51, 33, 7, 14]. Specifically, we follow their technical implementations (i.e., original settings) and utilize their source codes to reproduce three representative EIA (i.e., Attack-0–Attack-2) and three representative AIA (i.e., Attack-3–Attack-5) attacks for evaluating PrivTune.
@@
We adopt the PEFT method, specifically LoRA [13], in PrivTune, and employ the AdamW optimizer with a linear learning rate scheduler during fine-tuning, with an initial learning rate of $3\times 10^{-4}$. By default, we set $l=3$, $n=3$, $k=2$, and $\delta=0.6$ in our experiments. For evaluation, we use classification accuracy and ROUGE-L score to measure the Utility Performance (UP) on classification and generation tasks, respectively [18]. Additionally, Attack Success Rate (ASR) is used to assess the defense capability.